OpenStudy (anonymous):

write an equation in slope intercept form for the line that is parallel to the given line and that passes through the given point: 4y=x;(4,0)

OpenStudy (anonymous):

slope of the given line is m = 1/4 given point (4, 0) put in point-slope form: y - y1 = m(x - x1), where (x1, y1) is the given point. what do you have?

OpenStudy (anonymous):

so y-0=1/4(x-4)

OpenStudy (anonymous):

yep. but clean it up. your answer can also be expressed as: y=1/4(x-4) or y = 1/4*x -1

OpenStudy (anonymous):

o ok thanks I see how you did that. Thank you so much. I was very confused on this concept.